# assumeAssumption

### v1

@@ 4.2 Linear symmetries
Linear symmetries give rise to a conservation law in subgradient dynamics, as shown in [23]. We next recall it and compute the Hessian of the conserved quantity in some directions, since it appears in ˜2. Note that a conservation law in gradient dynamics was proposed earlier in [43, Proposition 5.1]
Assumption 3.
@@
Proposition 6.
Let ˜3 hold and $c(x)=\|C(x)-C(\overline{x})\|_{F}^{2}/4$. Then
@@
By [23, Corollary 1], for all $x\in\mathbb{R}^{n}$, $v\in\overline{\partial}f(x)$, and $\alpha\in\mathbb{R}$, we have $C(x+\alpha v)=C(x)+\alpha^{2}C(v)$ and thus
@@
Proposition 7.
Under ˜3, $\overline{x}\in\arg\mathrm{loc}\min\{|x|:x\in G\overline{x}\}\Longrightarrow C(\overline{x})=0$.
@@
where the second equality holds by ˜12. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
@@
The function $\tau$ is lower semicontinuous and continuous on its compact domain $\partial f(\overline{x})\times[-\epsilon/2,\epsilon/2]$ since $f$ is Lipschitz continuous near $\overline{x}$. Thus $\varphi\circ\gamma$ is real-valued on $\epsilon/2,\epsilon/2]$. Since 0 is a local maximum of $\varphi\circ\gamma$, we have $0\in\partial(\varphi\circ\gamma)(0)$ by Fermat’s rule [36, Theorem 10.1]. As the sum of the indicator of a closed convex set and a smooth function, $\tau$ is regular [36, Example 7.28]. [36, Corollary 10.11] then implies that for all $(v,t)\in\mathrm{dom}\tau$, one has
@@
Due to its bounded domain, the function $\tau(v,t)$ is level-bounded in $v$ locally uniformly in $t$ [36, 1.16 Definition]. We may thus apply [36, Theorem 10.13] on parametric subdifferentiation. For all $\overline{t}\in[-\epsilon/2,\epsilon/2]$, it yields
@@
In Theorem˜2, one can also take a maximal Bouligand subdifferential by ˜1. The converse of Theorem˜2 do not always hold.
Example 4.
The function $f(x)=|2x_{1}^{2}-x_{2}^{2}|$ is invariant under the natural action of $G=\{A\in\mathrm{GL}(2,\mathbb{R}):A^{T}DA=D\}$ where $D=\mathrm{diag}(2,-1)$. The only flat minimum is $(0,0)$ even though $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$.
@@
Observe that $f(x)=\langle x,Dx\rangle$. For all $A\in\mathrm{GL}(2,\mathbb{R})$, $f(Ax)=\langle Ax,DAx\rangle=\langle x,A^{T}DAx\rangle=\langle x,Dx\rangle=f(x)$. Thus $f$ is invariant under the natural action of $G$, whose Lie algebra is
@@
Thus $\mathrm{s}(\mathfrak{g})=\{0\}$ and $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$. By ˜10, $\mathrm{lip}f(x)=2\sqrt{4x_{1}^{2}+x_{2}^{2}}$ and thus $(0,0)$ is the only flat minimum by Corollary˜1. ∎
The converse of Theorem˜2 does hold in $\ell_{1}$-matrix factorization, as we show in Section˜4.3.
Corollary 6.
Under ˜3, if $\overline{x}$ is flat and $f$ is regular near $\overline{x}$, then there exists $\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}$ such that $C(\overline{v})=0$.
@@
This a consequence of ˜12, Corollary˜1, and Theorem˜2. ∎
Theorem˜2 can be generalized to higher orders, as follows.
Theorem 3.
@@
The proof naturally mirrors that of Theorem˜2. By assumption, $I_{n}$ is a local solution to
@@
where the second equality holds by ˜13. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
@@
following Lim [28, Section 4]. Since the constraint is qualified by gradient independence, the first-order optimality condition ensures the existence of $\overline{\lambda}\in\mathbb{R}$ such that
@@
As in the proof of Theorem˜2, we conclude that $C(\overline{v})=0$. ∎
Corollary 7.
@@
This a consequence of ˜13, Corollary˜2, and Theorem˜3. ∎
@@ 4.3 Matrix factorization
We seek to establish converse results to those developed in the previous section, in the case of matrix factorization, in a desire to characterize flat minima. Given $m,n,r\in\mathbb{N}^{*}$ and $M\in\mathbb{R}^{m\times n}$, the map $F:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\ni(X,Y)\mapsto XY-M\in\mathbb{R}^{m\times n}$ is invariant under the action of $\mathrm{GL}(r,\mathbb{R})$ on $\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ defined by $(X,Y,A)\mapsto(XA,A^{-1}Y)$. As shown in [23, Example 3], for any locally Lipschitz function $g:\mathbb{R}^{m\times n}\to\mathbb{R}$, $g\circ F$ then admits the conserved quantity $C(X,Y)=X^{T}X-YY^{T}$ (see [3, 15, 21, 31] for other derivations). We successively treat the cases where $g=\|\cdot\|_{1}$ and $g=\|\cdot\|_{F}$. We begin with a simple fact.
Fact 15.
@@
˜15 implies that for any diagonal matrix $D$ of order $n$ with nonnegative entries and $P\in\mathrm{O}(n)$, one has $DP=PD$ iff $D^{1/2}P=PD^{1/2}$. The optimal value in the lemma below is already known [37, Lemma 1], but the local-global property appears to be new. Also, the characterization of global minima via the balance condition on $X$ and $Y$ was stated in [12, Lemma 2.2], although the proof seems invalid.
Lemma 4.
$2\|M\|_{*}=\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $X^{T}X=YY^{T}$.
@@
If $XY=M$, the $U^{T}XYV=\Sigma$ and $2\|M\|_{*}=2\|\Sigma\|_{*}\leq\|U^{T}X\|_{F}^{2}+\|YV\|_{F}^{2}=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, with equality when $(X,Y)=(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$.
If $(X,Y)$ is a local minimum then $XY=M$ and $X^{T}X=YY^{T}$ by Proposition˜7. Conversely, consider some compact singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$. Then $V_{X}\Sigma_{X}^{2}V_{X}^{T}=U_{Y}\Sigma_{Y}^{2}U_{Y}^{T}$, so that $\Sigma_{X}=\Sigma_{Y}$ and $\Sigma_{X}^{2}=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}(V_{X}^{T}U_{Y})^{T}$. In the proof of [12, Lemma 2.2], one concludes that $V_{X}^{T}U_{Y}=I_{r}$, but that seems untrue (think of $\Sigma_{X}=I_{r}$). In fact, $\Sigma_{X}^{2}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}$ and thus $\Sigma_{X}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}$ by ˜15. Using a compact singular value decomposition $M=U\Sigma V^{T}$, this yields $XY=U_{X}\Sigma_{X}V_{X}^{T}U_{Y}\Sigma_{X}V_{Y}^{T}=U_{X}V_{X}^{T}U_{Y}\Sigma_{X}^{2}V_{Y}^{T}=U\Sigma V^{T}$, and so $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$. Thus $\|X\|_{F}^{2}+\|Y\|_{F}^{2}=2\|M\|_{*}$ and $(X,Y)$ is globally optimal. ∎
Using Lemma˜4, we can show that the converse of Theorem˜2 holds in $\ell_{1}$-matrix factorization.
Proposition 8.
@@
(i) $\Longrightarrow$ (ii) is obvious. (ii) $\Longrightarrow$ (iii) is due to Theorem˜2. (iii) $\Longrightarrow$ (i) Observe that $f(X,Y)=\|F(X,Y)\|_{1}$ where F(X,Y)=XYMF(X,Y)=XY-M. We thus compute $F(X,Y)(H,K)=XK+HY$ and $F^{\prime}(X,Y)^{*}(\Lambda)=(X^{T}\Lambda,\Lambda Y^{T})$. By ˜10,
@@
By assumption, there exists $\overline{\Lambda}\in[-1,1]^{m\times n}$ such that $(\overline{H},\overline{K})=(\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T},\overline{X}^{T}\overline{\Lambda})$. Hence $\overline{H}\hskip 1.13809pt\overline{K}=\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T}\overline{X}^{T}\overline{\Lambda}=\overline{\Lambda}\hskip 1.13809ptM^{T}\overline{\Lambda}$. Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$, by Lemma˜4,
@@
Given a matrix $A\in\mathbb{R}^{m\times n}$ with singular value decomposition $A=U\Sigma V^{T}$, let $S_{A}=\{v\in\mathbb{R}^{n}:|Av|=\|A\|_{2}|v|\}$ and $d_{A}=\dim S_{A}$, i.e., the multiplicity of the maximal singular value of $A$. Let $\widehat{U}$ (resp. $\widehat{V}$) denote the first $d_{A}$ columns of $U$ (resp. $V$), in other words, the maximal left (resp. right) singular vectors of $A$.
Fact 16.
@@
For all $v\in\mathbb{R}^{n}$, $|XYv|\leq\|XY\|_{2}|v|\leq\|X\|_{2}\|Y\|_{2}|v|$ and $|XYv|\leq\|X\|_{2}|Yv|\leq\|X\|_{2}\|Y\|_{2}|v|$. Thus $S_{XY}\supset S_{X}\cap YS_{Y}$ and, if $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$, then $S_{XY}\subset S_{X}\cap YS_{Y}$. Lastly, $S_{X}=\mathrm{Im}\hskip 0.85358pt\widehat{V}_{X}$ and $YS_{Y}=\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}$. ∎
@@
Lemma 5.
$2\|M\|_{2}=\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$ iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$.
@@
with equality exactly when $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$. This happens in particular when $(X,Y)=(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$. If $\|X\|_{2}\neq\|Y\|_{2}$, say $\|X\|_{2}<\|Y\|_{2}$, then $\|tX\|_{2}^{2}+\|t^{-1}Y\|_{2}^{2}<\|X\|_{2}^{2}+\|Y\|_{2}^{2}$ for all $t\in(1,\|Y\|_{2}^{1/2}\|X\|_{2}^{-1/2}]$. Suppose $\|XY\|_{2}<\|X\|_{2}\|Y\|_{2}$, then consider some singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$ and let $d_{X}$ denote the multiplicity of the maximal singular value of $X$. Consider $D_{t}=\mathrm{diag}(tI_{d_{X}},I_{r-d_{X}})$ where $t>0$. Let $X_{t}=XV_{X}D_{t}V_{X}^{T}$ and $Y_{t}=V_{X}D_{t}^{-1}V_{X}^{T}Y$. For all $t\in(0,1)$,
@@
by ˜16. ∎
In contrast to the multilayer case [33], it is possible to determine the maximal eigenvalue and eigenspace of the Hessian at any global minimum.
Proposition 9.
@@
Observe that $f(X,Y)=\|F(X,Y)\|_{F}^{2}/2$ where F(X,Y)=XYMF(X,Y)=XY-M. By ˜11, $\lambda_{1}(\nabla f(X,Y))=\|F^{\prime}(X,Y)\|_{2}^{2}$ so we compute $F^{\prime}(X,Y)(H,K)=HY+XK$. One has
@@
Let $\widehat{V}_{X}$ (resp. $\widehat{U}_{Y}$) denote the first $d_{X}$ (resp. $d_{Y}$) columns of $V_{X}$ (resp. $U_{Y}$). If $(H,K)\in E_{1}(\nabla^{2}((X,Y))$ and $\mathrm{Im}\hskip 0.85358ptH^{T}\cap\mathrm{Im}\hskip 0.85358ptK\neq\{0\}$, then $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}A^{T}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}B\neq\{0\}$, $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}\neq\{0\}$, and $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ by ˜16. ∎

### v2

@@ 4.2 Linear symmetries
Linear symmetries give rise to a conservation law in subgradient dynamics, as shown in [26]. We next recall it and compute the Hessian of the conserved quantity in some directions, since it appears in Assumption˜4.3. Note that a conservation law in gradient dynamics was proposed earlier in [51, Proposition 5.1]
Assumption 4.10.
@@
Proposition 4.11.
Let Assumption˜4.10 hold and $c(x)=\|C(x)-C(\overline{x})\|_{F}^{2}/4$ for some $\overline{x}\in\mathbb{R}^{n}$. Then
@@
By [26, Corollary 5.1], for all $x\in\mathbb{R}^{n}$, $v\in\overline{\partial}f(x)$, and $\alpha\in\mathbb{R}$, we have $C(x+\alpha v)=C(x)+\alpha^{2}C(v)$ and thus
@@
Proposition 4.12.
Under Assumption˜4.10, $\overline{x}\in\arg\mathrm{loc}\min\{|x|:x\in G\overline{x}\}\Longrightarrow C(\overline{x})=0$.
@@
where the second equality holds by Fact˜3.28. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
@@
The function $\tau$ is lower semicontinuous and continuous on its compact domain $\partial f(\overline{x})\times[-\epsilon/2,\epsilon/2]$ since $f$ is Lipschitz continuous near $\overline{x}$. Thus $\varphi\circ\gamma$ is real-valued on $\epsilon/2,\epsilon/2]$. Since 0 is a local maximum of $\varphi\circ\gamma$, we have $0\in\partial(\varphi\circ\gamma)(0)$ by Fermat’s rule [44, Theorem 10.1]. As the sum of the indicator of a closed convex set and a smooth function, $\tau$ is regular [44, Example 7.28]. [44, Corollary 10.11] then implies that for all $(v,t)\in\mathrm{dom}\tau$, one has
@@
Due to its bounded domain, the function $\tau(v,t)$ is level-bounded in $v$ locally uniformly in $t$ [44, 1.16 Definition]. We may thus apply [44, Theorem 10.13] on parametric subdifferentiation. For all $\overline{t}\in[-\epsilon/2,\epsilon/2]$, it yields
@@
In Theorem˜4.13, one can also take a maximal Bouligand subdifferential by Fact˜2.1. The converse of Theorem˜4.13 do not always hold.
Example 4.14.
The function $f(x)=|2x_{1}^{2}-x_{2}^{2}|$ is invariant under the natural action of $G=\{A\in\mathrm{GL}(2):A^{T}DA=D\}$ where $D=\mathrm{diag}(2,-1)$. The only flat minimum is $(0,0)$ even though $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$.
@@
Observe that $f(x)=\langle x,Dx\rangle$. For all $A\in\mathrm{GL}(2)$, $f(Ax)=\langle Ax,DAx\rangle=\langle x,A^{T}DAx\rangle=\langle x,Dx\rangle=f(x)$. Thus $f$ is invariant under the natural action of $G$, whose Lie algebra is
@@
Thus $\mathrm{s}(\mathfrak{g})=\{0\}$ and $C(x)=P_{\mathrm{s}(\mathfrak{g})}(xx^{T})=0$ for all $x\in\mathbb{R}^{2}$. By Fact˜3.25, $\mathrm{lip}f(x)=2\sqrt{4x_{1}^{2}+x_{2}^{2}}$ and thus $(0,0)$ is the only flat minimum by Corollary˜3.18. ∎
The converse of Theorem˜4.13 does hold in $\ell_{1}$-matrix factorization, as we show in Section˜4.3.
Corollary 4.15.
Under Assumption˜4.10, if $\overline{x}$ is flat and $f$ is regular near $\overline{x}$, then there exists $\overline{v}\in\arg\max\{|v|:v\in\partial f(\overline{x})\}$ such that $C(\overline{v})=0$.
@@
This a consequence of Fact˜3.28, Corollary˜3.18, and Theorem˜4.13. ∎
Theorem˜4.13 can be generalized to higher orders, as follows.
Theorem 4.16.
@@
The proof naturally mirrors that of Theorem˜4.13. By assumption, $I_{n}$ is a local solution to
@@
where the second equality holds by Fact˜3.29. Also, $\phi:\mathbb{R}^{n}\times G\rightarrow\overline{\mathbb{R}}$ and $\varphi:G\rightarrow\mathbb{R}_{-}$ are defined by
@@
following Lim [33, Section 4]. Since the constraint is qualified by gradient independence, the first-order optimality condition ensures the existence of $\overline{\lambda}\in\mathbb{R}$ such that
@@
As in the proof of Theorem˜4.13, we conclude that $C(\overline{v})=0$. ∎
Corollary 4.17.
@@
This a consequence of Fact˜3.29, Corollary˜3.19, and Theorem˜4.16. ∎
@@ 4.3 Matrix factorization
We seek to establish converse results to those developed in the previous section, in the case of matrix factorization, in a desire to characterize flat minima. Given $m,n,r\in\mathbb{N}^{*}$ and $M\in\mathbb{R}^{m\times n}$, the map
$\begin{array}[]{cccc}F:&\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}&\longrightarrow&\mathbb{R}\\ &(X,Y)&\longmapsto&XY-M\end{array}$
is invariant under the action of $\mathrm{GL}(r)$ on $\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ defined by
$(X,Y,A)\mapsto(XA,A^{-1}Y).$
As shown in [26, Example 5.3], for any locally Lipschitz function $g:\mathbb{R}^{m\times n}\to\mathbb{R}$, $g\circ F$ then admits the conserved quantity
$C(X,Y)=X^{T}X-YY^{T}$
(see [3, 17, 24, 36] for other derivations). A pair $(X,Y)$ is balanced if $X^{T}X=YY^{T}$ [21, Chapter 6]. We successively treat the cases where $g=\|\cdot\|_{1}$ and $g=\|\cdot\|_{F}^{2}/2$. We begin with a simple fact, which is a special case of [13, Proposition 1].
Fact 4.18.
@@
Fact˜4.18 implies that for any diagonal matrix $D$ of order $n$ with nonnegative entries and $P\in\mathrm{O}(n)$, one has $DP=PD$ iff $D^{1/2}P=PD^{1/2}$. In particular, given $A\succcurlyeq 0$, there is a unique $B\succcurlyeq 0$ such that $A=B^{2}$, called the square root of $A$. (Indeed if $B^{2}=\widetilde{B}^{2}$ with $B,\widetilde{B}\succcurlyeq 0$, then given some eigendecompositions $B=PDP^{T}$ and $\widetilde{B}=\widetilde{P}\widetilde{D}\widetilde{P}^{T}$, we have $D=\widetilde{D}$ and $\widetilde{P}^{T}PD^{2}=D^{2}\widetilde{P}^{T}P$. Thus $\widetilde{P}^{T}PD=D\widetilde{P}^{T}P$ and $B=\widetilde{B}$.)
The optimal value in the lemma below is already known [45, Lemma 1], but the local-global property appears to be new. Also, the characterization of global minima via the balance condition on $X$ and $Y$ was stated in [14, Lemma 2.2], although the proof seems invalid. A similar result constrained to a single orbit was obtained in [21, Theorem 3.7].
Lemma 4.19.
If $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, then $2\|M\|_{*}=\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $X^{T}X=YY^{T}$.
@@
If $XY=M$, then $U^{T}XYV=\Sigma$ and $2\|M\|_{*}=2\|\Sigma\|_{*}\leq\|U^{T}X\|_{F}^{2}+\|YV\|_{F}^{2}=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, with equality when $(X,Y)$ equals $(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$ up to some zero rows/columns. If $(X,Y)$ is a local minimum then $XY=M$ and $X^{T}X=YY^{T}$ by Proposition˜4.12. Conversely, consider some compact singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$. By Lemma˜4.20 below, $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$. Thus $\|X\|_{F}^{2}+\|Y\|_{F}^{2}=2\|M\|_{*}$ and $(X,Y)$ is globally optimal. ∎
The following result can also be deduced from [40, Lemma 3.1].
Lemma 4.20.
If $XY=M$ and $X^{T}X=YY^{T}$, then $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$ given some compact singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$
Proof.
Plugging in the decompositions yields $V_{X}\Sigma_{X}^{2}V_{X}^{T}=U_{Y}\Sigma_{Y}^{2}U_{Y}^{T}$, so that $\Sigma_{X}=\Sigma_{Y}$ and $\Sigma_{X}^{2}=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}(V_{X}^{T}U_{Y})^{T}$. In the proof of [14, Lemma 2.2], one concludes that $V_{X}^{T}U_{Y}=I_{r}$, but that seems untrue (think of $\Sigma_{X}=I_{r}$). In fact, $\Sigma_{X}^{2}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}^{2}$ and thus $\Sigma_{X}(V_{X}^{T}U_{Y})=(V_{X}^{T}U_{Y})\Sigma_{X}$ by Fact˜4.18. Using a compact singular value decomposition $M=U\Sigma V^{T}$, this yields $XY=U_{X}\Sigma_{X}V_{X}^{T}U_{Y}\Sigma_{X}V_{Y}^{T}=U_{X}V_{X}^{T}U_{Y}\Sigma_{X}^{2}V_{Y}^{T}=U\Sigma V^{T}$, and so $\Sigma_{X}=\Sigma_{Y}=\sqrt{\Sigma}$. ∎
Using Lemma˜4.19, we can show that the converse of Theorem˜4.13 holds in $\ell_{1}$-matrix factorization.
Proposition 4.21.
@@
(i) $\Longrightarrow$ (ii) is obvious. (ii) $\Longrightarrow$ (iii) is due to Theorem˜4.13. (iii) $\Longrightarrow$ (i) Observe that $f(X,Y)=\|F(X,Y)\|_{1}$ where F(X,Y)=XYMF(X,Y)=XY-M. We thus compute $F(X,Y)(H,K)=XK+HY$ and $F^{\prime}(X,Y)^{*}(\Lambda)=(X^{T}\Lambda,\Lambda Y^{T})$. By Fact˜3.25,
@@
By assumption, there exists $\overline{\Lambda}\in[-1,1]^{m\times n}$ such that $(\overline{H},\overline{K})=(\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T},\overline{X}^{T}\overline{\Lambda})$. Hence $\overline{H}\hskip 1.13809pt\overline{K}=\overline{\Lambda}\hskip 1.13809pt\overline{Y}^{T}\overline{X}^{T}\overline{\Lambda}=\overline{\Lambda}\hskip 1.13809ptM^{T}\overline{\Lambda}$. Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}$, by Lemma˜4.19,
@@
Given a real-valued matrix $A$ with a singular value decomposition $A=U_{A}\Sigma_{A}V_{A}^{T}$, let $S_{A}=\{v\in\mathbb{R}^{n}:|Av|=\|A\|_{2}|v|\}$ and $d_{A}=\dim S_{A}$, i.e., the multiplicity of the maximal singular value of $A$. Let $\widehat{U}_{A}$ (resp. $\widehat{V}_{A}$) denote the first $d_{A}$ columns of $U_{A}$ (resp. $V_{A}$), in other words, the maximal left (resp. right) singular vectors of $A$.
Fact 4.22.
@@
For all $v\in\mathbb{R}^{n}$, $|XYv|\leq\|XY\|_{2}|v|\leq\|X\|_{2}\|Y\|_{2}|v|$ and $|XYv|\leq\|X\|_{2}|Yv|\leq\|X\|_{2}\|Y\|_{2}|v|$. Thus $S_{XY}\supseteq S_{X}\cap YS_{Y}$ and, if $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$, then $S_{XY}\subseteq S_{X}\cap YS_{Y}$. Lastly, $S_{X}=\mathrm{Im}\hskip 0.85358pt\widehat{V}_{X}$ and $YS_{Y}=\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}$. ∎
@@
Lemma 4.23.
If $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, then $2\|M\|_{2}=\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$ and a feasible point $(X,Y)$ is a global minimum iff it is a local minimum iff $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$ iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$.
@@
with equality exactly when $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ and $\|X\|_{2}=\|Y\|_{2}$. This happens in particular when $(X,Y)$ equals to $(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$ up to some zero rows/columns. If $\|X\|_{2}\neq\|Y\|_{2}$, say $\|X\|_{2}<\|Y\|_{2}$, then $\|tX\|_{2}^{2}+\|t^{-1}Y\|_{2}^{2}<\|X\|_{2}^{2}+\|Y\|_{2}^{2}$ for all $t\in(1,\|Y\|_{2}^{1/2}\|X\|_{2}^{-1/2}]$. Suppose $\|XY\|_{2}<\|X\|_{2}\|Y\|_{2}$, then consider some singular value decompositions $X=U_{X}\Sigma_{X}V_{X}^{T}$ and $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$ and let $d_{X}$ denote the multiplicity of the maximal singular value of $X$. Consider $D_{t}=\mathrm{diag}(tI_{d_{X}},I_{r-d_{X}})$ where $t>0$. Let $X_{t}=XV_{X}D_{t}V_{X}^{T}$ and $Y_{t}=V_{X}D_{t}^{-1}V_{X}^{T}Y$. For all $t\in(0,1)$,
@@
by Fact˜4.22. ∎
Lemma˜4.19, Lemma˜4.20, and Lemma˜4.23 immediately imply the following.
Corollary 4.24.
$\arg\min\{\|X\|_{F}^{2}+\|Y\|_{F}^{2}:XY=M\}\subseteq\arg\min\{\|X\|_{2}^{2}+\|Y\|_{2}^{2}:XY=M\}$.
The inclusion can be strict in Corollary˜4.24, as we will see in Example˜4.29. In contrast to the multilayer case [38], it is possible to determine the maximal eigenvalue and eigenspace of the Hessian at any global minimum. In fact, shortly after submitting this work on arXiv, an expression for the maximal eigenvalue in the multilayer case was proposed using Kronecker products [30, Theorem 5], implying the expression below [30, Corollary 6].
Proposition 4.25.
@@
Observe that $f(X,Y)=\|F(X,Y)\|_{F}^{2}/2$ where F(X,Y)=XYMF(X,Y)=XY-M. By Fact˜3.26, $\lambda_{1}(\nabla f(X,Y))=\|F^{\prime}(X,Y)\|_{2}^{2}$ so we compute $F^{\prime}(X,Y)(H,K)=HY+XK$. One has
@@
Let $\widehat{V}_{X}$ (resp. $\widehat{U}_{Y}$) denote the first $d_{X}$ (resp. $d_{Y}$) columns of $V_{X}$ (resp. $U_{Y}$). If $(H,K)\in E_{1}(\nabla^{2}((X,Y))$ and $\mathrm{Im}\hskip 0.85358ptH^{T}\cap\mathrm{Im}\hskip 0.85358ptK\neq\{0\}$, then $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}A^{T}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}B\neq\{0\}$, $\mathrm{Im}\hskip 0.85358pt\widehat{U}_{Y}\cap\mathrm{Im}\hskip 0.85358pt\widehat{V}_{Y}\neq\{0\}$, and $\|XY\|_{2}=\|X\|_{2}\|Y\|_{2}$ by Fact˜4.22. ∎